# Bridging Language, Agent, and World Models for Reasoning and Planning

### v1

@@ 1 Introduction
Planning is a fundamental cognitive ability that involves generating sequences of actions and reasoning about future states [11, 28]. However, prior work has shown that large language models (LLMs) often struggle with planning tasks, especially those requiring multi-step reasoning or long-range dependencies [12, 26, 22]. At the same time, the crucial role of memory in supporting computational reasoning, planning, and other cognitive functions has been widely recognized across computing, computational neuroscience, and psychology [21, 19, 2, 10, 5]. Yet, while memory has been understood as useful to reasoning and adaptation, its structure and mechanisms in the natural language planning tasks remain largely underexplored in LLM-based multi-agent frameworks.
Inspired by working memory model in cognitive psychology [3], we propose EvoMem, a multi-agent framework that explicitly incorporates the dual-evolving memory into the planning process. The framework assembles three agents (Constraint Extractor, Verifier, and Actor) and two memory modules. Given a query, the Constraint Extractor identifies task-specific rules and constraints, establishing the Constraint Memory (CMem). This fixed memory anchors the Actor’s solution generation across iterations and updates only when a new query is introduced. The Verifier evaluates the Actor’s output and provides feedback for subsequent iterations, forming the Query-feedback Memory (QMem). This dynamic memory evolves within a single query, enabling solution refinement throughout the ongoing task. CMem and QMem are reset after each query.
EvoMem is conceptually aligned with Baddeley’s multi-component theory of working memory first proposed in 1974 [3] (Details shown in App. B ). CMem plays a role analogous to the phonological loop by maintaining all constraints in a stable, verbalized form across rounds. QMem functions similarly to the scratch pad, accumulating intermediate outputs, verifier feedback, and error signals into a coherent multi-round representation. The iterative Actor–Verifier process serves as a central executive that selectively attends to CMem, updates QMem, and directs the next reasoning step. This cognitive framework explains why combining long-lived constraints with dynamically updated, per-round information leads to a more consistent and effective planning process.
We evaluate EvoMem on the NaturalPlan benchmark, covering trip planning, calendar scheduling, and meeting planning tasks. Using Gemini-1.5-Pro [24] as the primary backbone, EvoMem outperforms strong baselines, achieving average gains of +11.17% in trip planning, +2.56% in calendar scheduling, and +3.76% in meeting planning. To assess model-agnostic performance, we further test on DeepSeek V3 [15] and GPT-4.1-mini [1], and observe consistent improvements across models.
@@ 2.1 Multi-Agent Design
Constraint Extractor  Constraints are the rules and limitations that a valid plan must satisfy. For instance, in trip planning tasks, constraints may involve the total trip duration, fixed-date events (e.g., weddings or conferences), city-specific stays, and available flights. Prior work [20] indicates the criticality of constraint identification for reliable plan verification. The constraint extractor derives task-specific constraints from the problem description, capturing the essential conditions that any candidate plan must satisfy. These constraints are then appended to the actor’s prompts, ensuring its reasoning remains grounded in the core requirements of the task.
@@ 3.1 Main Result
We conduct experiments on three datasets in NaturalPlan [35]: 1,000 instances each for Calendar Scheduling and Meeting Planning, and 1,600 instances for Trip Planning. For comparison, we adopt three baselines: (i) Zero-shot CoT [13], (ii) Self-Reflect [20], where the same model iteratively refines outputs through self-reflective feedback loops, and (iii) PlanGen [20], a strong method that achieves state-of-the-art results in multi-agent frameworks across these datasets. In EvoMem, we set the temperature of the constraint extractor to 0.1, the verifier to 0, and the actor to 0.7. We set the maximum number of iterations $T$ to 5.
@@ Appendix A Related Work
LLM Agents for Planning  Recent research has established that Large Language Models (LLMs) struggle to solve planning tasks directly from task descriptions [12, 22, 26]. This limitation has spurred the development of various natural language planning benchmarks and environments to study the problem comprehensively [35, 4, 33, 25]. In this work, we evaluate our approach on NaturalPlan [35], which includes trip planning, calendar scheduling, and meeting planning tasks.
In response to this challenge, a significant body of work has focused on creating LLM-based agent frameworks to enhance planning capabilities. Some approaches are tailored to specific domains [17, 31], Other methods [37, 32] focus on integrating external information to guide the planning process. Another line of research concentrates on optimizing the prompts themselves [7, 29].
The current SOTA LLM-based agent framework on natural language planning benchmarks is held by PlanGen [20]. While recent works continue to make progress, they have primarily focused on innovations within single-agent frameworks [14, 6].
@@
Memory for LLM Agents  The importance of memory in computational, planning and other cognitive tasks has been well established, e.g., by computing, computational neuroscience, cognition and psychology researchers [21, 19, 2, 9, 10, 5]. Memory mechanisms have become central to advancing LLM-based agents, as effective planning and reasoning often require retaining and reusing information beyond a single context window. Early efforts focused on memory utilization [8, 31, 18, 36, 16, 27], using short-term and long-term memory to support information storage and recall. More recent work has explored agentic memory [34, 23], where agents dynamically organize memories, establish new connections, and evolve their knowledge with experience. Another important direction is workflow memory [30], which captures and reuses common routines to guide future tasks and improve efficiency. However, none of the above methods investigate the effectiveness of memory mechanisms in natural language planning tasks. Our framework specifically addresses this gap.
@@ Appendix B Working Memory Model in Cognitive Psychology
The central executive is like the “manager" of your mind’s working memory. It controls your attention, decides what matters, and coordinates the other components. The visuospatial sketchpad is the “inner eye". It holds visual and spatial information—like imagining a map, remembering where objects are, or picturing a diagram—and becomes active whenever you visualize something in your head. The phonological loop is the “inner voice”. It keeps words or sounds in your mind for a short time. For example, you repeat a phone number to yourself so you do not forget it.
Figure 3: Working memory model proposed in 1974 [3].

### v2

@@ 1 Introduction
Planning is a fundamental cognitive ability that involves generating sequences of actions and reasoning about future states [jiao2024learning, wang2024planning]. However, prior work has shown that large language models (LLMs) often struggle with planning tasks, especially those requiring multi-step reasoning or long-range dependencies [kambhampati2024llms, valmeekam2023planning, shojaee2025illusion]. At the same time, the crucial role of memory in supporting computational reasoning, planning, and other cognitive functions has been widely recognized across computing, computational neuroscience, and psychology [schuurmans2023memory, momennejad2024memory, armeni2024transformer, gollwitzer2025psychology, brown2020stress]. Yet, while memory has been understood as useful to reasoning and adaptation, its structure and mechanisms in the natural language planning tasks remain largely underexplored in LLM-based multi-agent frameworks.
Inspired by working memory model in cognitive psychology [baddeley2012working], we propose EvoMem, a multi-agent framework that explicitly incorporates the dual-evolving memory into the planning process. The framework assembles three agents (Constraint Extractor, Verifier, and Actor) and two memory modules. Given a query, the Constraint Extractor identifies task-specific rules and constraints, establishing the Constraint Memory (CMem). This fixed memory anchors the Actor’s solution generation across iterations and updates only when a new query is introduced. The Verifier evaluates the Actor’s output and provides feedback for subsequent iterations, forming the Query-feedback Memory (QMem). This dynamic memory evolves within a single query, enabling solution refinement throughout the ongoing task. CMem and QMem are reset after each query.
EvoMem is conceptually aligned with Baddeley’s multi-component theory of working memory first proposed in 1974 [baddeley2012working] (Details shown in App. B ). CMem plays a role analogous to the phonological loop by maintaining all constraints in a stable, verbalized form across rounds. QMem functions similarly to the sketchpad, accumulating intermediate outputs, verifier feedback, and error signals into a coherent multi-round representation. The iterative Actor–Verifier process serves as a central executive that selectively attends to CMem, updates QMem, and directs the next reasoning step. This cognitive framework explains why combining long-lived constraints with dynamically updated, per-round information leads to a more consistent and effective planning process.
We evaluate EvoMem on the NaturalPlan benchmark, covering trip planning, calendar scheduling, and meeting planning tasks. Using Gemini-1.5-Pro [team2024gemini] as the primary backbone, EvoMem outperforms strong baselines, achieving average gains of +11.17% in trip planning, +2.56% in calendar scheduling, and +3.76% in meeting planning. To assess model-agnostic performance, we further test on DeepSeek V3 [liu2024deepseek] and GPT-4.1-mini [achiam2023gpt], and observe consistent improvements across models.
@@ 2.1 Multi-Agent Design
Constraint Extractor  Constraints are the rules and limitations that a valid plan must satisfy. For instance, in trip planning tasks, constraints may involve the total trip duration, fixed-date events (e.g., weddings or conferences), city-specific stays, and available flights. Prior work [parmar2025plangen] indicates the criticality of constraint identification for reliable plan verification. The constraint extractor derives task-specific constraints from the problem description, capturing the essential conditions that any candidate plan must satisfy. These constraints are then appended to the actor’s prompts, ensuring its reasoning remains grounded in the core requirements of the task.
@@ 3.1 Main Result
We conduct experiments on three datasets in NaturalPlan [zheng2024natural]: 1,000 instances each for Calendar Scheduling and Meeting Planning, and 1,600 instances for Trip Planning. For comparison, we adopt three baselines: (i) Zero-shot CoT [kojima2022large], (ii) Self-Reflect [parmar2025plangen], where the same model iteratively refines outputs through self-reflective feedback loops, and (iii) PlanGen [parmar2025plangen], a strong method that achieves state-of-the-art results in multi-agent frameworks across these datasets. In EvoMem, we set the temperature of the constraint extractor to 0.1, the verifier to 0, and the actor to 0.7. We set the maximum number of iterations $T$ to 5.
@@ Appendix A Related Work
LLM Agents for Planning  Recent research has established that Large Language Models (LLMs) struggle to solve planning tasks directly from task descriptions [kambhampati2024llms, shojaee2025illusion, valmeekam2023planning]. This limitation has spurred the development of various natural language planning benchmarks and environments to study the problem comprehensively [zheng2024natural, bohnet2024exploring, xie2024travelplanner, valmeekam2023planbench]. In this work, we evaluate our approach on NaturalPlan [zheng2024natural], which includes trip planning, calendar scheduling, and meeting planning tasks.
In response to this challenge, a significant body of work has focused on creating LLM-based agent frameworks to enhance planning capabilities. Some approaches are tailored to specific domains [liu2024tool, wu2024copilot], Other methods [zhu2024knowagent, xie2024human] focus on integrating external information to guide the planning process. Another line of research concentrates on optimizing the prompts themselves [chen2024reprompt, wang2023promptagent].
The current SOTA LLM-based agent framework on natural language planning benchmarks is held by PlanGen [parmar2025plangen]. While recent works continue to make progress, they have primarily focused on innovations within single-agent frameworks [lee2025evolving, chen2025sets].
@@
Memory for LLM Agents  The importance of memory in computational, planning and other cognitive tasks has been well established, e.g., by computing, computational neuroscience, cognition and psychology researchers [schuurmans2023memory, momennejad2024memory, armeni2024transformer, ClarkSuperSizing, gollwitzer2025psychology, brown2020stress]. Memory mechanisms have become central to advancing LLM-based agents, as effective planning and reasoning often require retaining and reusing information beyond a single context window. Early efforts focused on memory utilization [chhikara2025mem0, wu2024copilot, mei2024aios, zhong2024memorybank, liu2023think, wang2023enhancing], using short-term and long-term memory to support information storage and recall. More recent work has explored agentic memory [xu2025mem, tang2024enhancing], where agents dynamically organize memories, establish new connections, and evolve their knowledge with experience. Another important direction is workflow memory [wang2024agent], which captures and reuses common routines to guide future tasks and improve efficiency. However, none of the above methods investigate the effectiveness of memory mechanisms in natural language planning tasks. Our framework specifically addresses this gap.
@@ Appendix B Working Memory Model in Cognitive Psychology
The central executive is like the “manager" of your mind’s working memory. It controls your attention, decides what matters, and coordinates the other components. The visuo-spatial sketchpad is the “inner eye". It holds visual and spatial information—like imagining a map, remembering where objects are, or picturing a diagram—and becomes active whenever you visualize something in your head. The phonological loop is the “inner voice”. It keeps words or sounds in your mind for a short time. For example, you repeat a phone number to yourself so you do not forget it.
Figure 3: Working memory model proposed in 1974 [baddeley2012working].